OpenStudy (anonymous):

I have a graph, x over y. On top or x, is 2, 4, and 6. On bottom or y, is 10,20,30. I need to find the rate of change. Help please! I'm behind.

OpenStudy (anonymous):

rate of change is \[\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

so for the first pic (2,10) and (4,20) so \[\frac{20-10}{4-2} = \frac{10}{2} = 5\]

OpenStudy (anonymous):

now do the same for (4,20) and (6,30)

OpenStudy (anonymous):

You will see that the rate of change is 5.

OpenStudy (anonymous):

so that's it? It's just 5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Okay, thank you.